OpenStudy (anonymous):

consider the following series infinity ∑ n=1 x^n /5^n find the value of x for which the series converges answer in interval notation 2) find the sum of the series for those value of x. Write the formula in terms of x. Sum=

OpenStudy (anonymous):

How do I do it by steps?

OpenStudy (anonymous):

apply the ratio test, and force the ratio to be less than 1.

OpenStudy (anonymous):

Hi, how do I find the value of x for which the series converges answer in interval notation

OpenStudy (anonymous):

I'm still a bit confused...

OpenStudy (anonymous):

\[\left|\frac{a_{n+1}}{a_n}\right|<1\]

OpenStudy (anonymous):

Where \[a_n=\frac{x^n}{5^n}\]

OpenStudy (anonymous):

Thats how you do these problems in general. Although for this particular problem, you can use the fact that:\[\frac{x^n}{5^n}=\left(\frac{x}{5}\right)^n\]so your series is a geometric series, which only converges is the common ratio is less than 1.